OpenStudy (anonymous):

(x^3+9x^2-5x+11)/(x^2+2) solve using synthetic division if possible

OpenStudy (anonymous):

I know you cant use synthetic division b/c of the ^2 in (x^2+2) but i came up with an odd remainder using regular division

OpenStudy (anonymous):

you cannot use synthetic division for this because your denominator has degree 2

OpenStudy (anonymous):

So the quotient is \[\frac{s^3+9x^2-5x+11}{x^2+2}=x+9-\frac{7x+7}{x^2+2}\]